# LTM: Scalable and Black-box Similarity-based Test Suite Minimization based on Language Models

### v3

@@ 4.3.4 RQ2 results
Figure 3 shows that UniXcoder/Cosine converges faster to a higher FDR ($0.84$ across projects), compared to ATM/Combined ($0.80$ across projects) and ATM/TreeEditDistance ($0.81$ across projects). This may be explained by the fact that (1) UniXcoder utilizes both code comments and AST information from the source code, whereas ATM ignores code comments, and (2) UniXcoder was pre-trained on various code understanding tasks to help the model learn the contextual information for each token, and thus the meaning of each code element (e.g., method, variable names, and string characters) in the input test code. In contrast, ATM uses tree-based similarity based on ASTs, which lack information regarding the syntactic context and certain details of the test code.
For some projects, ATM achieved higher FDR than LTM. However, a Fisher’s Exact Test suggests that, for these projects, there is no significant difference between the FDR results of ATM and LTM. Nevertheless, a possible explanation for such performance difference could be that the percentage of test code exceeding the token limits of language models on these projects (5.72%, 3.21%, 3.25%, and 7.25% for $Chart$, $Compress$, $Collections$, and $Math$, respectively) is higher than the average of 2.98% across projects, and that the truncated parts are important for distinguishing test cases. Another possible explanation could be due to variability across test cases in some projects (e.g, $Cli$ and $Time$), regarding their low-level code elements such as complex string manipulation and domain-specific character strings, rather than structural information. This makes tree-based similarity better at capturing these differences, by comparing AST node labels, than using the Cosine Similarity of embeddings.
@@
Achieved Saving in Testing Time. The average time saving rate of the best configuration of LTM ranges from $18.13\%$ to $58.03\%$ across projects, with a mean of $41.72\%$ across projects. We observe that this saving rate is slightly higher than ATM/Combined ($40.03\%$) and ATM/TreeEditDistance ($41.02\%$).
@@
The results of quadratic regression confirmed that the MT and the number of test cases, for both LTM and ATM, follow a quadratic relationship as the coefficients on the quadratic term were statistically significant with $\alpha=0.01$. This is, as expected, due to the fact that (1) similarity values are measured between test case pairs, which increases quadratically with the number of test cases, and (2) the search space for GA increases rapidly with the number of test cases. The $R^{2}$ values for the regression model of LTM and ATM/Combined were $98.89\%$ and $89.51\%$, respectively. This indicates that a high percentage of the variation in MT can be explained by the number of test cases. However, the $R^{2}$ value for ATM/TreeEditDistance was only $50.36\%$, indicating that a much smaller proportion of the variation in MT is be explained by the number of test cases in this case. This is because the preparation time of ATM/TreeEditDistance, which is part of its MT, was also largely influenced by AST sizes. More discussions will be provided in the following section.
@@
Preparation Time. We found that LTM is much more efficient, in terms of preparation time across project versions (mean = $0.23$ min and median = $0.14$ min), compared to ATM/Combined (mean = $3.15$ min and median = $0.74$ min) and ATM/TreeEditDistance (mean = $35.31$ min and median = $6.10$ min). Moreover, the preparation time of LTM takes only $35.75\%$ of the MT per version, that is much less than ATM/Combined ($71.22\%$) and ATM/TreeEditDistance ($93.74\%$). Quadratic regression shows that the coefficient on the quadratic term for LTM ($1.241e^{-06}$) is one order of magnitude smaller than the coefficients for ATM/Combined ($6.634e^{-05}$) and two orders of magnitude smaller than ATM/TreeEditDistance ($4.3680e^{-04}$), with all coefficients statistically significant with $\alpha=0.01$. This indicates that, on larger industry projects with much larger test suites, the absolute calculation time difference between LTM and ATM is expected to be extremely large, thus making LTM a better option regarding scalability in preparation time.
The $R^{2}$ values for the regression model of LTM and ATM/Combined are $97.70\%$ and $87.86\%$, respectively. This indicates that more than $87\%$ of the variation in preparation time can be explained by the number of test cases. However, the $R^{2}$ value for ATM/TreeEditDistance is $49.96\%$, which is mainly due to its preparation time as it is highly influenced by test code length. For instance, as shown in Figure 4, there are project versions with nearly $4k$ test cases from the $Time$ project, which have less preparation time than project versions with around $3k$ test case from the $Math$ project. This is due to that fact that the average test code for the $Math$ project is longer than that for the $Time$ project, resulting in its AST tree size being twice as long. Therefore, although the test suite size of the $Math$ project is smaller, its preparation time is significantly longer than that of the $Time$ project. In contrast, the preparation time for LTM is not affected by the length of test code, which further supports our claims regarding its scalability.
@@
Note that the search time not only depends on the number of test cases per version but is also impacted by the number of generations required by the GA to converge. For instance, though the number of test cases per version for the $Math$ project ($2,078$) is less than that for the $Time$ project ($3,918$), GA needed more generations to converge for $Math$, thus resulting in longer search time. Test cases of the $Math$ project are more diverse than that of the $Time$ project, with a higher standard deviation of similarity values ($0.13$ compared to $0.09$), thus leading to more generations and longer convergence time for the GA search.
@@ 4.4 Discussion
Our results show that LTM runs up to two orders of magnitude faster than the SOTA approach (i.e., ATM) while achieving a statistically significantly higher fault detection rate. We should note that, according to the average MT across projects, LTM runs five times faster than ATM/Combined and $44$ times faster than ATM/TreeEditDistance. These differences tend to increase significantly as the test suite size grows. For example, for the $Math$ project with $2078$ test cases, LTM runs $6$ and $101$ times faster than ATM/Combined and ATM/TreeEditDistance, respectively. This suggests that LTM, while achieving higher effectiveness, can save enormous time and resources for minimizing test suites of the much larger software systems and test suites encountered in practice. The preparation phase plays a crucial role in the scalability of LTM. Since we employed pre-trained language models to extract vector-based test method embeddings and used a highly optimized function to compute similarity between them, we managed to achieve much shorter preparation times compared to ATM. In addition, unlike ATM, the preparation time of LTM is not impacted by test code length, which further supports our claims about its scalability. Furthermore, since we used the test code as input without preprocessing it and converting it into other formats, such as ASTs for ATM, we saved substantial memory space as well.
Unlike coverage-based techniques, the size of a project (i.e., source lines of code) has no impact on the minimization time of black-box TSM approaches, including LTM and ATM, as minimization time mostly depends on the number of test cases per project version. In addition, we optimize the search process by changing the data structure for input similarity values and accelerating the fitness calculation, which reduced the search time by $272.90$ times. The $Closure$ project—the largest by far, comprising $179$ KLoC and $174$ versions, with $7,308$ test cases per version—took only $17.80$ min in terms of MT and achieved a FDR of $0.79$, for a 50% minimization budget. Note that for ATM, $Closure$ was far too large to even run our experiments, thus further illustrating the scalability gains with LTM.
@@
LTM used the source code of test cases without preprocessing as input to the language models and yielded better results than using preprocessed code, as shown in our preliminary study. However, some preprocessing techniques are still worth investigating, such as normalizing variable identifiers, since they might not have an impact on the data flow in the source code. Therefore, using such techniques may potentially help language models generate embeddings that are more generalizable and suitable for code similarity calculation across projects, thus improving the fault detection power of LTM.
@@ 5 Threats to validity
Internal Validity Internal threats to validity are concerned with the ability to draw conclusions from our experimental results. To make fair comparisons and draw valid conclusions, we conducted our experiments using the same experimental design and dataset used to evaluate ATM [8]. We evaluated the alternative LTM configurations and compared the best configuration to the two best configurations of ATM. For minimizing test suites, we used the same fitness function and parameter settings as ATM, but we further optimized GA by utilizing a more efficient data structure for storing the input similarity values and accelerating the fitness calculation. We compared the performance using optimized GA for both LTM and ATM. However, ATM used preprocessed code as input, whereas LTM used source code without preprocessing. We did not consider using the full code for ATM due to its scalability issues. Another threat is related to the token length limit of language models, which might truncate the source code of some test cases [49], thus resulting in information loss that might in turn negatively impact the performance of LTM. However, we observed that, for each version, the percentage of test methods with the token length exceeding the limit is only $2.98\%$ across projects, which can be considered negligible. Nevertheless, some projects have a relatively higher percentage of test methods exceeding that limit (e.g., $7.25\%$ for $Math$), which might explain the slightly lower FDR of LTM compared to ATM on these projects. Future research should further investigate the use of additional language models with longer token length limits, though they might greatly impact efficiency.
@@ 6 Conclusion
In this paper, we proposed LTM, a scalable and black-box similarity-based test suite minimization (TSM) approach based on pre-trained language models. We investigated five alternative language models—CodeBERT, GraphCodeBERT, UniXcoder, StarEncoder, and CodeLlama—to extract test method embeddings and two similarity measures (i.e., Cosine Similarity and Euclidean Distance) for similarity calculation based on these embeddings. We employed an optimized version of Genetic Algorithm (GA) by utilizing an efficient data structure for handling similarity values to improve fitness calculation and finding optimal subsets of test suites. We assessed the performance of alternative LTM configurations on a large dataset consisting of $17$ Java projects with a total of $835$ versions collected from Defects4J. We used the Fault Detection Rate (FDR), Minimization Time (MT), and Time Saving Rate (TSR) as evaluation metrics to assess the effectiveness and efficiency of LTM. We identified the best configuration of LTM and compared it to ATM, a recent black-box TSM approach that achieved better trade-off in terms of effectiveness and efficiency than the alternative SOTA TSM approach (FAST-R). Results indicate that the best LTM configuration (i.e., UniXcoder/Cosine) outperformed the two best configurations of ATM by achieving significantly higher FDR results ($0.84$ versus $0.81$, on average) and more importantly, running much faster ($0.82$ min versus $4.06$ min, on average) than ATM, in terms of both preparation time (up to two orders of magnitude faster) and search time (one order of magnitude faster), which is particularly important on much larger industrial systems and test suites, while achieving a saving of $41.72\%$ in test execution time for a $50\%$ test minimization budget.

### v4

@@ 4.3.4 RQ2 results
Figure 3 shows that UniXcoder/Cosine converges faster to a higher FDR ($0.84$ across projects), compared to ATM/Combined ($0.80$ across projects) and ATM/TreeEditDistance ($0.81$ across projects). This may be explained by the fact that (1) UniXcoder can handle full code (i.e., the source code of test cases without preprocessing), whereas ATM uses preprocessed code; and (2) UniXcoder was pre-trained on various code understanding tasks, enabling the model to learn contextual information for each token and the meaning of each code element (e.g., method, variable names, and string characters) in the input test code. In contrast, ATM relies on tree-based similarity based on ASTs, which lack information regarding the syntactic context and certain details of the test code.
For some projects, ATM achieved higher FDR than LTM. However, a Fisher’s Exact Test suggests that, for these projects, there is no significant difference between the FDR results of ATM and LTM. Nevertheless, a possible explanation for such performance difference could be that the percentage of test code exceeding the token limits of language models on these projects (5.72%, 3.21%, 3.25%, and 7.25% for $Chart$, $Compress$, $Collections$, and $Math$, respectively) is higher than the average of 2.98% across projects, and that the truncated parts are important for distinguishing test cases. Another possible explanation could be due to variability across test cases in some projects (e.g., $Cli$ and $Time$), regarding their low-level code elements such as complex string manipulation and domain-specific character strings, rather than structural information. This makes tree-based similarity better at capturing these differences, by comparing AST node labels, than using the Cosine Similarity of embeddings.
@@
Achieved Saving in Testing Time. The average time saving rate of the best configuration of LTM ranges from $18.13\%$ to $58.03\%$ across projects, with a mean of $41.72\%$ across projects. We observe that this time saving rate is slightly higher than, but not significantly different from, that of ATM/Combined ($40.30\%$) and ATM/TreeEditDistance ($41.02\%$), as shown by the Wilcoxon signed-rank test.
Comparing LTM and ATM using preprocessed code. As discussed in Section 4.2.1, ATM cannot handle full code due to scalability issues. However, the preprocessing step might remove useful information from the test code and thus impact the effectiveness of ATM. Therefore, in order to explain the LTM FDR results when compared to ATM, we want to determine how the use of full code and the LTM test method embeddings individually affect FDR. For this, we conducted an additional experiment to assess LTM with UniXcoder/Cosine, the best LTM configuration, using preprocessed code, on all $16$ projects. Tables II and IV report the FDR and TSR results, respectively, for this configuration using preprocessed code (UniXcoder-p) and the two best configurations of ATM (using GA with combined and tree edit distance similarity measures), for a $50\%$ minimization budget. We observe that, using the same input as ATM (preprocessed test code), LTM with UniXcoder/Cosine still yields a higher average FDR ($0.83$) than ATM ($0.81$). Still, it is lower than what we achieved using LTM with full code ($0.84$). The average TSR result for LTM with preprocessed code is however slightly lower than that of ATM ($38.84\%$ versus $41.02\%$), though the difference is not statistically significant, as resulted by the Wilcoxon signed-rank test. These results therefore suggest that the small improvements in FDR observed with LTM are due to both LTM’s test method embeddings and its capacity to handle full code in a scalable way. TSR remains similar for LTM and ATM, with or without preprocessing.
@@
The results of quadratic regression confirmed that the MT and the number of test cases, for both LTM and ATM, follow a quadratic relationship as the coefficients on the quadratic term were statistically significant with $\alpha=0.01$. This is, as expected, due to the fact that (1) similarity values are measured between test case pairs, which increases quadratically with the number of test cases, and (2) the search space for GA increases rapidly with the number of test cases. The $R^{2}$ values for the regression model of LTM and ATM/Combined were $98.89\%$ and $89.51\%$, respectively. This indicates that a high percentage of the variation in MT can be explained by the number of test cases. However, the $R^{2}$ value for ATM/TreeEditDistance was only $50.36\%$, indicating that a much smaller proportion of the variation in MT can be explained by the number of test cases in this case. This is because the preparation time of ATM/TreeEditDistance, which is part of its MT, was also largely influenced by AST sizes. More discussions will be provided in the following section.
@@
Preparation Time. We found that LTM is much more efficient, in terms of preparation time across project versions (mean = $0.23$ min and median = $0.14$ min), compared to ATM/Combined (mean = $3.15$ min and median = $0.74$ min) and ATM/TreeEditDistance (mean = $35.31$ min and median = $6.10$ min). Moreover, the preparation time of LTM takes only $35.75\%$ of the MT per version, which is much less than ATM/Combined ($71.22\%$) and ATM/TreeEditDistance ($93.74\%$). Quadratic regression shows that the coefficient on the quadratic term for LTM ($1.241e^{-06}$) is one order of magnitude smaller than the coefficients for ATM/Combined ($6.634e^{-05}$) and two orders of magnitude smaller than ATM/TreeEditDistance ($4.3680e^{-04}$), with all coefficients statistically significant with $\alpha=0.01$. This indicates that, on larger industry projects with much larger test suites, the absolute calculation time difference between LTM and ATM is expected to be extremely large, thus making LTM a better option regarding scalability in preparation time.
The $R^{2}$ values for the regression model of LTM and ATM/Combined are $97.70\%$ and $87.86\%$, respectively. This indicates that more than $87\%$ of the variation in preparation time can be explained by the number of test cases. However, the $R^{2}$ value for ATM/TreeEditDistance is $49.96\%$, which is mainly due to its preparation time as it is highly influenced by test code length. For instance, as shown in Figure 4, there are project versions with nearly $4k$ test cases from the $Time$ project, which have less preparation time than project versions with around $3k$ test case from the $Math$ project. This is due to the fact that the average test code for the $Math$ project is longer than that for the $Time$ project, resulting in its AST tree size being twice as long. Therefore, although the test suite size of the $Math$ project is smaller, its preparation time is significantly longer than that of the $Time$ project. In contrast, the preparation time for LTM is not affected by the length of the test code, which further supports our claims regarding its scalability.
@@
Note that the search time not only depends on the number of test cases per version but is also impacted by the number of generations required by the GA to converge. For instance, as shown in Figure 4, though the number of test cases per version for the $Math$ project ($2,078$) is less than that for the $Time$ project ($3,918$), GA needed more generations to converge for $Math$, thus resulting in longer search time. Test cases of the $Math$ project are more diverse than that of the $Time$ project, with a higher standard deviation of similarity values ($0.13$ compared to $0.09$), thus leading to more generations and longer convergence time for the GA search.
@@ 4.4 Discussion
Our results show that LTM runs up to two orders of magnitude faster than the SOTA approach (i.e., ATM) while achieving a statistically significantly higher fault detection rate. We should note that, according to the average MT across projects, LTM runs five times faster than ATM/Combined and $44$ times faster than ATM/TreeEditDistance. These differences tend to increase significantly as the test suite size grows. For example, for the $Math$ project with $2078$ test cases, LTM runs $6$ and $102$ times faster than ATM/Combined and ATM/TreeEditDistance, respectively. This suggests that LTM, while achieving higher effectiveness, can save enormous time and resources for minimizing test suites of the much larger software systems and test suites encountered in practice. The preparation phase plays a crucial role in the scalability of LTM. Since we employed pre-trained language models to extract vector-based test method embeddings and used a highly optimized function to compute similarity between them, we managed to achieve much shorter preparation times compared to ATM. In addition, since we used the test code as input without preprocessing it and converting it into other formats, such as ASTs for ATM, we saved substantial memory space as well.
Unlike coverage-based techniques, the size of a project (i.e., source lines of code) has no impact on the minimization time of black-box TSM approaches, including LTM and ATM, as minimization time mostly depends on the number of test cases per project version. In addition, we optimize the search process by changing the data structure for input similarity values and accelerating the fitness calculation, which reduced the search time by $190.48$ times. The $Closure$ project—the largest by far, comprising $179$ KLoC and $174$ versions, with $7,308$ test cases per version—took only $17.80$ min in terms of MT and achieved a FDR of $0.79$, for a 50% minimization budget. Note that for ATM, $Closure$ was far too large to even run our experiments, thus further illustrating the scalability gains with LTM.
@@
We recommend that LTM be used with the source code of test cases without preprocessing, since it yielded better results than when using preprocessed code, as reported in Section 4.3.1 and also in Section 4.3.4. However, some aspects of preprocessing, such as normalizing variable identifiers, may be beneficial, since they do not have an impact on the data flow in the source code. Therefore, using such techniques may potentially help language models generate embeddings that are more generalizable and suitable for code similarity calculation across projects, thus improving the fault detection power of LTM.
@@ 5 Threats to validity
Internal Validity Internal threats to validity are concerned with the ability to draw conclusions from our experimental results. To make fair comparisons and draw valid conclusions, we conducted our experiments using the same experimental design and dataset used to evaluate ATM [8]. We evaluated the alternative LTM configurations and compared the best configuration to the two best configurations of ATM. For minimizing test suites, we used the same fitness function and parameter settings as ATM, but we further optimized GA by utilizing a more efficient data structure for storing the input similarity values and accelerating the fitness calculation. We compared the performance using optimized GA for both LTM and ATM.
ATM used preprocessed code as input because of its very long preparation time. In contrast, LTM used full source code, raising the question of whether differences in FDR between LTM and ATM are due to differences in test representation (test method embeddings or ASTs) or code input (full code or preprocessed code). To answer that question, we conducted an additional experiment to assess the best configuration of LTM (i.e., UniXcoder/Cosine) using preprocessed code, which showed that LTM with preprocessed code still fares better than ATM in terms of FDR.
Another threat is related to the token length limit of language models, which might truncate the source code of some test cases [49], thus resulting in information loss that might in turn negatively impact the performance of LTM. However, we observed that, for each version, the percentage of test methods with the token length exceeding the limit is only $2.98\%$ across projects, which can be considered negligible. Nevertheless, some projects have a relatively higher percentage of test methods exceeding that limit (e.g., $7.25\%$ for $Math$), which might explain the slightly lower FDR of LTM compared to ATM on these projects. Future research should further investigate the use of additional language models with longer token length limits, though they might greatly impact efficiency.
@@ 6 Conclusion
In this paper, we proposed LTM, a scalable and black-box similarity-based test suite minimization (TSM) approach based on pre-trained language models. We investigated five alternative language models—CodeBERT, GraphCodeBERT, UniXcoder, StarEncoder, and CodeLlama—to extract test method embeddings and two similarity measures (i.e., Cosine Similarity and Euclidean Distance) for similarity calculation based on these embeddings. We employed an optimized version of the Genetic Algorithm (GA) by utilizing an efficient data structure for handling similarity values to improve fitness calculation and find optimal subsets of test suites. We assessed the performance of alternative LTM configurations on a large dataset consisting of $17$ Java projects with a total of $835$ versions collected from Defects4J. We used the Fault Detection Rate (FDR), Minimization Time (MT), and Time Saving Rate (TSR) as evaluation metrics to assess the effectiveness and efficiency of LTM. We identified the best configuration of LTM and compared it to ATM, a recent black-box TSM approach that achieved better trade-off in terms of effectiveness and efficiency than the alternative SOTA TSM approach (FAST-R). Results indicate that the best LTM configuration (i.e., UniXcoder/Cosine) outperformed the two best configurations of ATM by achieving significantly higher FDR results ($0.84$ versus $0.81$, on average) and more importantly, running much faster ($0.82$ min versus $4.06$ min, on average) than ATM, in terms of both preparation time (up to two orders of magnitude faster) and search time (up to one order of magnitude faster), which is particularly important on much larger industrial systems and test suites, while achieving a saving of $41.72\%$ in test execution time for a $50\%$ test minimization budget.